solve 4+ln(27x) = 25-2lnx
Answer by lwsshak3(11628)

solve 4+ln(27x) = 25-2lnx
ln(27x)+2lnx=21
ln[(27x)(x^2)]=21
ln(27x^3)=21
convert to exponential form
e^21=27x^3
take cube root of both sides
3x=e^7
x=e^7/3